OpenStudy (anonymous):

Which of the following are solutions to the quadratic equation below? Check all that apply. x^2 - 4x - 32 = 0 a. 8 b. 2 c. -1 d. -4 e. 1 f. 6

OpenStudy (anonymous):

ans is 8, -4....

jimthompson5910 (jim_thompson5910):

x^2-4x-32=0 (x-8)(x+4) = 0 x-8 = 0 or x+4 = 0 x = 8 or x = -4 So the answers are A and D